### v1

# SkyAI Sim: An Open-Source Simulation of UAV Aerial Imaging from Satellite Data

Abstract
Capturing real-world aerial images for vision-based navigation (VBN) is challenging due to limited availability and conditions that make it nearly impossible to access all desired images from any location. The complexity increases when multiple locations are involved. The state of the art solutions, such as flying a UAV (Unmanned Aerial Vehicle) to take pictures or using existing research databases, have significant limitations. SkyAI Sim offers a compelling alternative by simulating a UAV to capture bird’s-eye view satellite images at zero-yaw with real-world visible-band specifications. This open-source tool allows users to specify the bounding box (top-left and bottom-right) coordinates of any region on a map. Without the need to physically fly a drone, the virtual Python UAV performs a raster search to capture satellite images using the Google Maps Static API. Users can define parameters such as flight altitude, aspect ratio and diagonal field of view of the camera, and the overlap between consecutive images. SkyAI Sim’s capabilities range from capturing a few low-altitude images for basic applications to generating extensive datasets of entire cities for complex tasks like deep learning. This versatility makes SkyAI a valuable tool for not only VBN, but also other applications including environmental monitoring, construction, and city management. The open-source nature of the tool also allows for extending the raster search to other missions. A dataset of Memphis, TN has been provided along with this simulator, partially generated using SkyAI and, also includes data from a 3D world generation package for comparison.
Keywords Vision-Based Navigation  $\cdot$ UAV  $\cdot$ Satellite Image  $\cdot$ Aerial Imaging  $\cdot$ Simulation  $\cdot$ Dataset  $\cdot$ Google Map API
1 Introduction
Figure 1: Image captured at coordinates (35.128039, -89.799163) at above ground level altitude ($AGL$) of 126 m. Image captured (a) by a UAV and simulated (b-d) with a field of view ($FOV$) of 78.8 degrees and aspect ratio ($AR$) of 4:3. Figures b-d represent simulation using three different maptypes on Google Map Static API.
1.1 Application
Aerial imagery is essential for many applications, such as construction, city administration, environmental monitoring, and VBN. In artifitial intelligence (AI) and machine learning (ML), these images provide training data for models, aiding in accurate analysis and decision-making. In VBN, they help autonomous systems, like drones, identify obstacles and landmarks [24]. Environmental monitoring relies on aerial imagery to assess changes in land use, track deforestation, and monitor wildlife habitats [16]. In construction, they enhance project management by offering site overviews, improving resource allocation, and ensuring safety [17]. For city management, they facilitate urban planning, infrastructure maintenance, and disaster response [11]. Additionally, aerial images are key in AI/ML for developing object detection, classification, and segmentation algorithms [15, 18, 22].
1.2 Related Work
Acquiring real-world aerial images is challenging due to constraints. The limited availability of images is a primary issue, as it is often impractical to obtain the required aerial views from specific locations. This difficulty is compounded when there is a need to cover multiple locations on Earth. Deploying Unmanned Aerial Vehicles (UAVs), have inherent limitations. For instance, flying a UAV to capture aerial images involves logistical complexities, significant time investment, and financial costs. Moreover, UAV operations are subject to regulatory restrictions, such as airspace regulations and flight permissions, as well as environmental factors like weather conditions that can impede flight missions.
While there are many valuable datasets, they have limitations. Every dataset is designed for a specific applications such as segmentation, object detection, and more. Reference [12] gives a comprehensive collection of many of these datasets. There are real-world aerial images such as FloodNet [20, 21]. Floodnet includes 2343 samples of flooded areas and provides a comprehensive dataset for classification, semantic segmentation, and visual question answering. On the other hand, many datasets are derived from satellite data. The University1652-Baseline dataset [26, 25] provides both drone-based and satellite-view geo-localization datasets. While these datasets are highly valuable for their specific tasks, they are limited by the parameter selection of the authors. They cannot replace the benefit of choosing preferred locations and altitudes as a real drone mission would.
Although datasets are useful, they may not cover the exact scenarios needed for comprehensive testing or algorithm development. Simulators offer a controlled environment to generate large amounts of synthetic data, customize conditions, and adapt to new models without the constraints of physical data collection. They also provide repeatability, safety, and cost-effectiveness, addressing issues like incomplete data or challenging environments that might be difficult to replicate with real-world datasets.
SkyAI Sim  [^1] https://github.com/JacobsSensorLab/SkyAI-Sim addresses these challenges by offering a versatile tool to capture geo-tagged simulated aerial images. It emulates a UAV to capture bird’s-eye view satellite images that adhere to real-world visible-band specifications, with the drone facing north at a zero yaw as represented in Figure 1. A bird’s-eye view is a 90-degree tilt, means the drone’s camera is looking straight down, orthogonal to the ground. However, this limitation is not significant for most applications as many applications predominantly use bird’s-eye view images. With SkyAI, users can simulate various missions using the Google Maps Static API in Python, by specifying the coordinates of any region on a map. Furthermore, if the users have the tools to collect their own drone-view images, SkyAI expands the capabilities of the University1652 dataset to any desired location. This enables them to perform tasks based on both drone-view and satellite-view images.
1.3 Google Maps Static API
The Google Maps Static API allows embedding a static map image on a webpage without needing JavaScript or dynamic content loading. The API generates a map based on specified settings and provides it as an image through a standard HTTP request with URL parameters. This simplifies the integration of maps into webpages, offering customization options including map center, zoom level, size, and type, making it versatile for adding geographic visualizations without dynamic interactions [3]. The API uses the Web Mercator Projection, a cylindrical map projection that preserves angles and shapes but distorts size and distance, especially near the poles [5, 9]. Understanding this projection aids in accurately mapping dimensions from the API to the real world.
2 Materials and Methods
In this section, a step-by-step guide will be provided, starting with the theoretical calculations required to define various missions. We will also delve into the architecture of the program, offering an explanation for users to not only utilize, but also advance the program to more sophisticated levels.
2.1 Mission Definition
To capture a single image using a drone, it is required to collect all necessary materials and resources similar to a real drone flight, such as altitude and camera settings. Once a single image is captured, we set a solid foundation for more advanced missions. For example, we can program the drone to take multiple images at different locations or complete a raster mission by covering a specified area, given the bounding box coordinates.
2.1.1 Single Image Capture
In order to use Sky-AI Sim, there are some flight and camera parameters that have to be known. To capture a single image, these parameters include geodetic latitude $\phi$ and longitude $\lambda$, above ground level altitude in meters ($AGL$), the diagonal field of view in degrees ($FOV$), and the imaging aspect ratio $AR$ of the camera. It is essential to remember that in this version, the camera tilt is 90 degrees and the drone yaw is 0. In satellite simulation, the image pixel size ($res_{x,y}$) and zoom level ($z$) respectfully resemble the $FOV$ and $AGL$ in the real world. Digital dynamic maps are made possible by loading a small portion of the globe map data at a time, allowing for rapid panning and zooming. Google Maps uses the Mercator map projection, and the world is represented as $256\times 256$ pixel tiles. The zoom levels on the map, ranging from 0 to 23 or higher, are represented by a grid of these tiles organized in a pyramidal pattern. The scale of the map is determined by each zoom level, which is a numerical scalar. For instance, the zoom ranges from 0 to 22 in our real-world simulation if we use the satellite format. Tiles show generic information such as continents and oceans at lower zoom levels, resembling flying at a very high altitude. On the other hand, more detailed views of streets and buildings are displayed at lower altitudes and bigger zoom levels. A grid of $2^{z}\times 2^{z}$ determines the number of tiles at each zoom level ($z$). We may use the formula $C_{e}/2^{z}$ level to find the real size of a single tile on a given zoom level, where $C_{e}$ is the circumference of the earth (40,075,017 meters). Hence, the number of meters per tile side is obtained [8].
Table 1: Guide to understand zoom level. The "# Tiles" column shows the number of tiles required to cover the Earth at each zoom level, useful for calculating storage needs for pre-generated tiles. "Tile width" column indicates the map width in degrees of longitude per square tile at that zoom level. The "m / pixels" column lists the meters per pixel at the Equator for 256-pixel-wide tiles, based on an Earth radius of 6372.7982 km; these values should be adjusted by the cosine of the latitude for other regions [10, 8].
Level Tile width (°) m / pixel (on Equator)
m / tile slide
(on Equator)
Examples of areas to represent
0 360 156 543 40,075,017 whole world
1 180 78 272 20,037,504
2 90 39 136 10,018,752 subcontinental area
3 45 19 568 5,009,376 largest country
4 22.5 9 784 2,504,688
5 11.25 4 892 1,252,344 large African country
6 5.625 2 446 626,172 large European country
7 2.813 1 223 313,086 small country, US state
8 1.406 611.496 156,543 US large national park
9 0.703 305.748 78,272 large metropolitan area
10 0.352 152.874 39,136 metropolitan area
11 0.176 76.437 19,568 city
12 0.088 38.219 9,784 town, or city district
13 0.044 19.109 4,892 village, or suburb
14 0.022 9.555 2,446 residential area, airport
15 0.011 4.777 1,223 small road
16 0.005 2.389 611 street
17 0.003 1.194 306 block, park
18 0.001 0.5972 153 some buildings, trees
19 0.0005 0.2986 76 local highway, crossing details
20 0.00025 0.1493 38 mid-sized building
21 0.000125 0.0746 19 a house
22 0.0000625 0.0373 10 a single apartment
The process of capturing a map from the Google Maps Static API [3], involves understanding the relationship between the zoom level, image pixel size, and geolocation bounds. The API allows one to specify the zoom level of the map. The zoom level determines the scale of the map and, consequently, the resolution of the captured image. Higher zoom levels provide more detail, while lower zoom levels cover larger areas with less detail. As follows, this section explains how the simulation parameters $AGL,FOV,AR$ are converted to parameters required to capture satellite data including $z$ and $res_{x,y}$  [^2] The functions introduced in ”Single Image Capture” section endnotes are located in the ”src/utils” folder within the ”geo_helper” package..
  • Map Dimensions in Meters:
    To calculate the image diagonal in meters ($d_{diag}$) given $AGL$ and $FOV$, use:
    $d_{diag}=2\cdot AGL\cdot\tan\left(\beta\cdot FOV/2\right)$ (1)
    where $\beta$ is a scaler to convert degrees to radians ($\pi/180$). Knowing $d_{diag}$, the width of the map along the x and y-axes ($d_{x}$, $d_{y}$) in meters can be calculated as follows  [^3] The calculations of land size in meters can be performed using the ”get_map_dim_m” function.:
    $\begin{split}d_{x}=&d_{diag}\cdot\sin\left(\tan^{-1}(AR)\right)\\ d_{y}=&d_{diag}\cdot\cos\left(\tan^{-1}(AR)\right)\end{split}$ (2)
  • Determine Geolocation Bounds from Metrics
    Knowing the dimensions in meters, with the benefits of the geopy package [13], the bounding box coordinates will be determined  [^4] The process of determining geolocation bounds from metric information is performed by the ”calc_bbox_m” function.. This step can be skipped if the bounding box is already known.
  • Determine Zoom Level and Pixel Size from Bounding Box
    The goal of this section is to find the zoom level that maximizes the pixel-wise resolution for the specified area. This involves: 1. using the geolocation bounds to determine the map’s span in degrees; 2. calculating the necessary zoom level to fit the entire span within the image dimensions while providing the highest resolution. We can achieve the goal of this section, by inversing the function explained in the next step.
  • Determine Bounding Box from Zoom Level and Pixel Size
    The Google Maps JavaScript API [1] The Google Maps JavaScript API’s getBounds() function provides the latitude and longitude boundaries of the visible map area based on the central coordinates, zoom level, and map size in pixels ($res_{x,y}$, also referred to as the resolution of the map). To simulate a geographic area from a central point, zoom level, and map dimensions, we calculate the bounding box coordinates using Mercator projection principles. The following steps detail this calculation process. 555”calc_bbox_api” function implements bounding box calculation in Python.
    • Center Point Offset: Offsetting the center point from the center of a web Mercator projection tile $(128,128)$ according to latitude and longitude ($x_{c}$ and $y_{c}$) [5]:
      $x_{c}$ $=128+\alpha\cdot\beta\cdot\lambda$ (3)
      $y_{c}$ $=128+\frac{\alpha}{2}ln\frac{1+sin(\beta\cdot\phi)}{1-sin(\beta\cdot\phi)}$
      $\alpha$ is a constant to scale the longitude from radians to pixels to represents the number of pixels per radians ($256/2\pi$).
    • Bounding Box in Mercator Tile: The map’s resolution in pixels is scaled to the tile in Mercator projection, with the help of a scale factor called pixel size. This parameter is calculated from the zoom level ($z$):
      $\Delta p=2^{-z}$ (4)
      Knowing the scaler ($\Delta p$), it is also possible to obtain the scaled width and height of the region of interest using:
      $res^{\prime}_{x,y}=res_{x,y}\cdot\Delta p$ (5)
      where $res^{\prime}_{x,y}$ is the resolution in pixels mapped from the real world to the web mercator projection. The bounding box pixels are expnaded by half the width from the central point.
    • Convert to coordinates: Convert pixel values to coordinates using the inverse of Eqn. 3 ($pnt$ is the top-left ($TL$) or bottom-right ($BR$) point.):
      $\lambda_{pnt}$ $=(x_{pnt}-128)/(\alpha\cdot\beta)$ (6)
      $\phi_{pnt}$ $=sin^{-1}(tanh((y_{pnt}-128)/(\alpha\cdot\beta)))$
    Knowing how to calculate the bounding box (getting the top left and bottom right latitude and longitude from the pixel size and zoom level of the geolocation of the center of the image), it is possible to reverse the process. In the reverse process, the bounding box is known. Hence, the center is known. The maximum possible $res_{x,y}$ accessible from the static API is $640\times 640$. With the bounds, finding $res^{\prime}_{x,y}$ is even easier: $res^{\prime}_{x}=x_{BR}-x_{TL}$. The same applies to the y-axis. Reversing Eqn. 4:
    $z=\lfloor-log_{2}(max(res^{\prime}_{x},res^{\prime}_{y})/640)\rfloor$ (7)
    Finding the maximum of the resolution’s width and height ensures the image fits within the selected zoom level. Using Eqn. 4 and Eqn. 5, the actual resolution of the image can be measured. 666 ”get_zoom_from_bounds” function finds a zoom level that can maximize the resolution of the captured image by the API.
  • Get Static Map Image
    Finally, once the optimal zoom level is determined, the API is used to capture the image. This is done by setting the map’s center, zoom level, and image pixel size on the API’s URL. For realistic images, the map type is set to "satellite" by default. However, the program is not limited to capturing realistic images and different map types can be requested [4]. Using "satellite" data, the zoom level can range between 0 and 22 as mentioned in Table 1. It’s important to note that the zoom level limits vary across different map types. The directory for where to save the captured data has to be determined. A flowchart of the process is available in Figure 3. An API key should be generated to access the API  [^7] The ”init_static_map” takes in the converted information and returns a URL response in the output. It is important to note that this function removes all the street names and text labels from the downloaded image. This can be undone by removing the ”style” parameter in the function. API-key can be fed to the program directly, or specified in a file named ”hidden_file.py” in the source code directory. This key should not be shared publicly and should stay secure. ,  [^8] The URL output can be downloaded locally using ”get_static_map_image” function. The ”retry” parameter determines the number of retries if the download fails..
2.1.2 Raster Image Mission
A logical extension of a single image capture, can define various missions. The main purpose of SkyAI Sim is to capture raster images of a specified area. In this mission, the objective is to complete a zero-tilt raster image capture, as shown in Fig. 2. The capture begins once the bounding box coordinates of the mission’s rectangular region (the map) are set. The coordinates are converted to UTM x and y to facilitate metric calculations and maintain the mission within a rectangular boundary. To ensure the mission remains within one or two UTM zones, it is advisable to restrict its scope to a practical distance. The first image is centered at the top left coordinates of the map. Zoom and image resolution will be calculated from the previous section by inputting the camera $FOV$, $AR$, and the drone’s $AGL$. By default the images are not overlapped and the distance between two consecutive images is the width of the image. Overlap can be accounted for, however. In this situation, the distance will be $d^{\prime}=d(1-overlap)$ where $overlap$ is the proportion of shared width/height to the total and has a value between 0 and 1. The mission ends when the bottom right coordinate is captured.
Figure 2: Raster mission visualization within UTM zone 16, provided bounding box coordinates.
Figure 3: The flowchart to process the map in downloading picture/s, using various missions. Each block includes the name of the function used available in the "src/utils" folder within the "geo_helper" package in parantheses. The "Download Raster" Section refers to the flowchart in Figure 4.
Figure 4: Raster search image capture flowchart.
2.2 Machine Learning PreProcessing
Since machine learning requires a substantial amount of data, raster missions are valuable for providing this dataset. However, not all raster samples may contain useful features for deep learning. Additionally, as the altitude decreases, the likelihood of obtaining featureless images increases. Therefore, it becomes increasingly important to filter out images that lack significant features. A few examples in Figure 5 demonstrate this.
Figure 5: Satellite samples. The image with a red border indicates no significant feature.
Shannon entropy can be used to clean data by identifying and removing samples with insignificant features. It measures the uncertainty or randomness in data, indicating the complexity of an image. High entropy suggests more complexity, while low entropy indicates order [2]. To calculate entropy: 1. Convert the image to grayscale; 2. Calculate pixel intensity histogram; 3. Normalize the histogram to get the probability distribution of pixel intensities. 4. Calculate the Shannon entropy using:
$H=-\sum_{i=0}^{255}p_{i}\log_{2}(p_{i})$ (8)
where $p_{i}$ is the probability of the pixel intensity $i$. Entropy values help identify samples with fewer significant features. Their values are stored in metadata, and the ones below a certain threshold can be removed later. By looking at the samples in Figure 5 using the roadmap mode (as illustrated in Figure 6) it is easier to detect the ones with fewer significant features.
Figure 6: The same images as in Figure 5 but in roadmap type. In roadmap, it is easier to detect lack of significant features as marked in red. The entropy of each image is marked on top of it.
3 Usage and Results
3.1 SkyAI Dataset
Using SkyAI-Sim, some sample data has been generated and published in this paper available in the "dataset" folder. "Memphis_single_sample" is an example of using the "download_single" program."random_samples" folder includes sample data listed in the "sample_coords.txt". This data is collected using "download_from_list" program. The main dataset includes the raster images of the Memphis Agricenter Area in Figure 7 referred to as Agricenter dataset. The images are collected in the rectangular bound between top left coordinates of $(35.16,-89.90)$ and bottom right coordinates of $(35.115,-89.823)$ using the "download_raster" program. The simulation has been done using a UAV based camera with a $FOV$ of 78.8 degrees and $AR$ of 4:3. The flight takes place at a fixed $AGL$ of 120 m. There are two sets of 1806 images each in "satellite" and "roadmap" modes without any overlap. According to the publish limits from the Google Map side, we could not publish a bigger dataset with this paper, however, a bigger dataset from Memphis area helped with the project described in [14]. This data is placed in "Memphis" folder under two subfolders "satellite_0" and "roadmap_0". This dataset is specifically designed to be used for machine learning tasks. For ease of access, the images are geotagged; a metadata table is also generated named "meta_data.csv". This file contains 6 columns: 1. "img_names" has the name of all images available. 2. "columns" and 3. "rows" consequently refer to the column and row indices of the raster search. 4. "Lat", 5. "Lon" also refer to the central coordinates of each image. 6. "Alt" for this dataset stores the zoom level value for each image. Some corresponding ML processings will be explained as follows.
Figure 7: Maps representing the area where the raster data has been collected in "satellite" mode on the left and "roadmap" on the
3.2 Supplementary Dataset
As previously noted, there are restrictions on publishing data from the API, making the SkyAI datasets themselves, insufficient for machine learning purposes. However, a more extensive dataset can be obtained from publicly available road data. In light of this, we are offering a larger second dataset that covers the city of Memphis. This additional dataset will serve as a valuable resource and can be used for comparison with the data generated by SkyAI Sim in various machine learning tasks.
NVIG, or the Night Vision Image Generator, is used to generate terrain for for capturing with sensors created in NVIPM or Night Vision Integrated Performance Model. The OpenStreetMap dataset [6] was used to generate a representation of the Memphis area depicted in Figure 7 for NVIG. The primary data that OpenStreetMap provides us are accurate road coordinates. By downloading all the road data within the bounding box coordinates using QGIS program [19] tools, we get shape files which are then parsed to extract the vectors of road coordinates. The most approachable data to extract after the roads are the trees. By creating a threshold function to mask out the regions of satellite imagery of the bounded Memphis area, we can fill the masked regions with random coordinates of varying densities to simulate the presence of trees.
With the road and tree features extracted as series and individual coordinates respectively, the remaining task was generating the appropriate road and tree files for NVIG to render the features properly. After the tools were developed to automate this, the addition of roads and trees to the NVIG terrain simulation was quick and efficient. However, some caveats that resulted from this technique were artifacts in the road that manifested as sharp and unnatural spikes in the terrain of which the cause has not been discovered. Another problem comes from the way roads are generated in NVIG. The roads are not fitted to the terrain; instead, they are formed by points and try to fit to the terrain curvature using a very broad slope parameter. A workaround has been to elevate the height of the road so that it is offset some distance above the ground. As most of the imagery is aerial, this does not pose a great issue; however, it would be problematic if imagery were to be collected from a ground level. Additionally, the road widths are rarely specified in OpenStreetMap. This forces us to use an arbitrary road width when it is not provided. As a result, the road width is not a reliable metric for feature extraction with the NVIG terrain.
Generating buildings was also a consideration for the NVIG Agricenter terrain dataset; however, the complexity and effort to implement automated building placement in NVIG deterred us from taking that route. The end result is a medium fidelity 3D representation of the Agricenter terrain over which custom simulated sensors can collect data. Despite the lack of buildings and the occasional road artifacts, the availability of coordinate-accurate roads in a 3D space is powerful in and of itself.
The bounding box for this data has a top left coordinate of (35.218752, -90.075361), and bottom right (35.064913, -89.730661). The images are taken in an $AGL$ of 300 m with horizontal and vertical $FOV$ of 75 and make a total of 3350 images. This additional dataset will enhance our ability to generate road data and will serve as a key point of comparison with the data from SkyAI Sim. By comparing the two datasets, we can better assess their strengths and limitations, ultimately improving the accuracy of our machine learning models specifically for road extraction tasks. Figure 8 illustrates one sample of this dataset.
(a) 35.113862, -89.801008
(b) 35.111531, -89.808043
(c) 35.118524, -89.843216
Figure 8: Sample road data extracted from OpenStreetMap dataset simulated with the NVIPM model. Central coordinates are provided for each sample.
3.3 Usage
To start using the API, you first need a project with a billing account on Google Maps API. When integrating the Google Maps Static API into SkyAI-Sim, it’s crucial to understand the API’s limitations and Google’s policies on content usage  [^9] https://about.google/brand-resource-center/products-and-services/geo-guidelines/#required-attribution/. The code architecture is as follows.
Figure 9: Code Structure.
The project consists of two main folders: 1. "src" has all the python codes in it. 2. "datasets" is the predefined location to store images to, or load the stored data from. There are more details about this folder and its content in the dataset section. The focus of this section is on the "src" folder and the code structure.
  • skyai.data: The data package has three main data modules in it:
    1. 1.
      ImageData: The abstract module to manage image like data.
    2. 2.
      VBN: The module inherited from ImageData with more specific features for VBN tasks. This module can be used to load the NVIG dataset.
    3. 3.
      SkyAI: This is the main module to load the SkyAI data. It is inherited from VBN module and has its specific attributes and methods to handle GoogleMap static API.
  • skyai.utils: The utility package for managing more generic tasks:
    • consts: Stores all constant values.
    • config: Manages input parsed arguments.
    • geo_helper: Includes all helper functions for geolocation based calculations. Different parts of this module have been explicitly explained in the Missions section.
    • img_helper: Consists of all helper functions required to mange images in general.
    • io_helper: Helper functions that help with inputting and outputing data is in this package. It connects the code to the local directories and the terminal output.
    • preprocess: Additional image preprocessing tools are provided in this package.
  • skyai.notebooks: There are two main notebooks available in this directory They both take care of processing of the datasets that are currently $d$atacleaner" notebook helps with looking at the entropies of random samples and preparing the histogram and analysis on the data cleaner threshold selection. "preprocess" notebook eases the visualization of applying various preprocessing tools on random sample images from the downloaded data.
3.3.1 Input Formats for the Argument Parser
The argument parser for running the script, accepts multiple parameters as input.
  1. 1.
    coords: The coordinates is entered as a string. It can have different formats:
    1. (a)
      Address to a file: If the input is a string referring to a file, the file should contain coordinates and $AGL$ values in the order of 1. $\phi$, 2. $\lambda$, 3. $AGL$ in meters, per line, separated by a space as indicated below. When processing a single download file, only the first row will be considered.
      $\phi_{1}$ $\lambda_{1}$ $AGL_{1}$
      $\phi_{2}$ $\lambda_{2}$ $AGL_{2}$
    2. (b)
      Formatted String Input: Alternatively, the input can be provided as a string in one of the following formats:
      • Raster Coordinates: Top-left and bottom-right coordinates for raster download. The input format should be:
        "$\phi_{TL}$_$\lambda_{TL}$_$\phi_{BR}$_$\lambda_{BR}$_$AGL$"
        For example:
        "35.22_-90.07_35.06_-89.73_400"
        In the case of a single download, the top-left coordinate will be used as the central coordinate, and the bottom-right coordinate will be discarded.
      • Central Coordinates and AGL: The input can also be provided as the central coordinates for a single image capture:
        "$\phi$_$\lambda$_$AGL$"
  2. 2.
    fov: Diagonal field of view in degrees. The default is 78.8.
  3. 3.
    aspect-ratio: The aspect ratio width and height separated by a space, e.g., 4 3 for a 4:3 ratio which is the default value.
  4. 4.
    map_type: The type of static map for realistic view is set to "satellite" by default. It can be adjusted to "roadmap" or "terrain" as well.
  5. 5.
    dataset: A string to specify which dataset is going to be used.
  6. 6.
    The default is "SkyAI". It can be switched to "VBN" to access and process the Supplementary data.
  7. 7.
    data_dir: The directory to read/write the data from/to.
  8. 8.
    vmargin: The vertical margin is an optional but crucial parameter. Google permits publishing images with watermarks, but these can bias data. To avoid this, we capture images with an extra vertical margin—20% total, split as 10% at the top and 10% at the bottom. This margin is later cropped out during processing by the "preprocess_image" method in the Google Maps module.
  9. 9.
    img_size: After capturing the data, they can get resized to any desired resolution specified by this parameter. It takes three parameters:1. width, height, and number of channels (1 for grayscale and 3 for RGB) separated by a space.
  10. 10.
    overlap: This parameter is set to 0 by default and is only effective for the raster download program. It is the proportion of shared width/height to the total and has a value between 0 and 1.
  11. 11.
    batch_size: Batch size is only an effective parameter if post processing machine learning dataset is used. It is set to 8 by default.
  12. 12.
    seed: An initial value used by a random number generator to produce a sequence of numbers, ensuring the same sequence can be reproduced in future runs. The default is 2024 and it can be adjusted if necessary.
These parameters all have their default values in the "config.py" file located in the "src/utils" directory. If other parameters than the default values is desired, specifically for the coordinates, they can be adjusted in the command line terminal, as a bash file, or as a "config.json" file. The command line has the highest priority. If no values are provided there or in a bash script, the "config.json" file will be used. If neither is set, the defaults in "config.py" will apply.
3.4 Post Processing
To cleanup the data as an optional step, a threshold entropy is defined. The threshold is selected according to the distribution of entropies in all images, along with experimental observations provided in Figure 6. Specifically in the Agricenter dataset, the images below an entropy of 2.1 lack significant details. To confirm this value’s reliability and ensure it doesn’t filter out substantial valuable data, a barplot and histogram of entropies for all available roadmap images are shown in Figures 10 and 11. This ensures most data is above the threshold, making it safe to discard images below $2.1$. SkyAI performs data cleaning based on the map type. If the map type is "satellite" or anything other than "roadmap", it first checks for "roadmap" data in the data directory. If available, images are filtered according to "roadmap" images with an entropy above the default threshold (2.1). If no roadmap data exists, images are filtered based on their original map type.
Figure 10: Bar plot showing the entropies of all roadmap images, with most images having an entropy above the 2.1 threshold.
Figure 11: Histogram of image entropies with the red line marking the 2.1 threshold. As shown, a significant portion of data falls below this threshold, but these samples mostly consist of unremarkable features, like green areas with trees and farms. Discarding them is ideal, as the majority of valuable data remains safely above the threshold.
After the data is cleaned using the "cleanup_data" method, the "config_dnn" method can be called to prepare the dataset for a machine learning task. In this process, the labels are the central latitude and longitude of each image, which are then converted to UTM and normalized using a standard scaler. The data is then split into three sets: train, validation, and test. At this stage, the cleaned data is transferred to a Keras dataset [23], where it is preprocessed and divided into batches. Data augmentation is then applied, including random rotation, flip, and zoom by default. Additional options like random contrast and brightness adjustments are also available if needed.
4 Discussion
SkyAI Sim opens the door to numerous applications, but it does have its limitations. Some of these are inherited from the API it relies on. A notable limitation is the inability to determine the exact date of the imagery captured through the Google Maps Static API. Google does not provide this information via the API, and the dates can vary widely depending on the region and zoom level. This is due to Google Maps using a mix of satellite, aerial, and Street View imagery from different times. Since the Google Maps Static API does not provide the date of image capture, this limitation, while not critical for most of SkyAI Sim’s potential applications, is still an important area for future improvement. One possible solution is to integrate additional data sources that offer time-stamped satellite imagery. For instance, APIs from services like Google Earth Engine [7] could be leveraged to retrieve historical satellite images with precise dates. This would enhance the accuracy and utility of static map images in applications where temporal information is crucial.
Another limitation is the restriction on the number of images that can be freely published. If SkyAI Sim is used to generate and publish a specific dataset, no more than 5,000 images can be released without obtaining permission from Google. Additionally, free access to data retrieval is not unlimited; monthly restrictions apply to the amount of data that can be retrieved for free. However, these limitations are manageable, as obtaining permission or making additional payments can easily resolve the issue.
Lastly, the current setup, which captures images at zero yaw with a 90-degree tilt, limits the range of available angles. However, this constraint presents an opportunity for future enhancement. Future versions could incorporate more advanced algorithms and utilize APIs beyond the static Google Maps API to dynamically adjust both tilt and yaw. This would expand the range of angles available, offering greater flexibility and more comprehensive data capture capabilities.
5 Conclusion
SkyAI Sim’s functionality allows users to customize various parameters, including flight altitude, camera aspect ratio, diagonal field of view, and image overlap. This flexibility supports a wide range of applications, from capturing low-altitude images for simple tasks to generating comprehensive datasets for complex projects such as deep learning. In addition to its versatility, SkyAI Sim’s open-source nature encourages further development and adaptation to other missions beyond its initial design. The tool’s ability to produce extensive datasets without the need for physical UAV deployment represents a significant advancement in dataset generation for VBN algorithm developments. SkyAI Sim presents a compelling solution to the challenges of aerial image capture, offering a powerful and flexible tool for a variety of applications, and opening new avenues for research and development in VBN and beyond.